OpenStudy (anonymous):

Find six distinct points on the circle with center (2,3) and radius 5

OpenStudy (anonymous):

equation of a circle with center (x0, y0) and radius R is: (x - x0)² + (y - y0)² = R²

OpenStudy (anonymous):

if we plug this values, we get: (x - 2)² + (y - 3)² = 25

OpenStudy (anonymous):

now make combinations of x and y so that this equation is satisfied

OpenStudy (anonymous):

plz make 1 combination so i can make another

OpenStudy (anonymous):

is there any method for making combination

OpenStudy (anonymous):

if i set x = 2, then: (y-3)² = 25 y-3 = 5 y = 8, so point (2,8) is in the circle

OpenStudy (anonymous):

we have another one, since \[\sqrt{25} = \pm 5\]

OpenStudy (anonymous):

so y-3 = -5 y = -2, so point (2,-2) is also a point in the circle

OpenStudy (anonymous):

now, set y = 3 and find more two values for x, after that we already have four points

OpenStudy (anonymous):

we can put any value of x and find y ? or is there any rule for assuming the value of x ?

OpenStudy (anonymous):

no, you can't

OpenStudy (anonymous):

we have a range of possible values for x and y

OpenStudy (anonymous):

how i know what are these range

OpenStudy (anonymous):

the center of the circle is (2,3) and radius 5, so x is inside range x = [2-5,2+5] and y = [3-5,3+5]

OpenStudy (anonymous):

oh i see thanks a lot thank u so much

OpenStudy (anonymous):

but you must choose one of these point, plug in the equation to find the assotiated values

OpenStudy (anonymous):

because this range that i gave you, is a ractangle, not a circle, what defines a circle it's equation

OpenStudy (anonymous):

the range which u gave me is correct ? i can chose that one ?